OpenStudy (anonymous):

how to solve: 5/2x-8/9=1/18-1/3x and also determine the values of the variable that make a denominator 0

OpenStudy (bradely):

5/2x-8/9=1/18-1/3x 5/2x+1/3x =1/18+8/9 17/6x=17/18 Solve this to find x Source: http://www.mathskey.com/question2answer/